# Learning User Preferences Through Interaction for Long-Term Collaboration

### v1

@@ 8 Conclusion
Our work provides a foundation for multi-session collaboration, with many promising directions for future research. The user simulators in MultiSessionCollab can be extended to have more complex preferences, such as context-dependent preferences that vary across domains, or evolving preferences that develop over time, enabling the evaluation of more sophisticated adaptation abilities. As user simulation advances, simulators can also express preferences through more diverse means, for example through implicit behavioral cues. Finally, improving memory management abilities and addressing ensuring agents do not forget useful information as they update memories remains an important direction.
@@ 9 Limitations
We use Llama-3.3-70B-Instruct as our user simulator. While it is a highly capable model, user simulation requires reliably following complex instructions, such as consistently enforcing preferences throughout interactions, and our evaluation metrics depend on this behavior. We found that the simulator struggled to exhibit expected behaviors during more complex tasks such as coding, which we consequently excluded from our benchmark. Additionally, user simulators may not be realistic or fully capture all kinds of human behavior, for example real users can express preferences implicitly or adapt to the agent over time. Although we conduct a human user study to validate our findings across realistic tasks and real users, we acknowledge these limitations remain and merit further investigation.
@@ Appendix A User Interaction Preferences
In human-AI collaborative settings, users often have specific interaction preferences that describe how they expect agents to behave. When agents adhere to these preferences, users may learn more effectively, communicate more efficiently, or experience smoother, more personalized interactions (shi2025models; sweller1988cognitive). To ensure our benchmark captures realistic preferences, we draw inspiration from studies in psychology, cognitive science, and human-computer interaction. We manually curate these preferences to be well-suited for our collaborative problem-solving setting. Table LABEL:tab:interaction_preferences provides the complete taxonomy of interaction preferences used in the MultiSessionCollab benchmark.
@@
You prefer agent responses to avoid any unnecessary narration and get directly to the point. Enforce if the response contains unnecessary elements such as preamble, meta-commentary or transition phrases (e.g. ’That’s a great question!’ or ’Let me walk you through this…’). You prefer agent responses to be narrative, conversational, and engaging. The response should acknowledge previous context (e.g. ’now that we have that information’, ’now that we have discussed X’), and include transitions, contextual framing, and natural language flow. Enforce if the response is overly terse or choppy (e.g. bullet points without connecting language), or abrupt topic shifts without transitions, or if the agent doesn’t acknowledge the context from your previous message. gudykunst1988culture
@@
You prefer the agent responses to be respectful, considerate, and friendly. Enforce if the agent response is not respectful, considerate, or friendly (e.g. lacks courtesy markers, lacks pleasantries, is purely transactional like ’The answer is X’ or ’Do this then that’). You prefer the agent responses to be blunt without pleasantries, apologetic language, or courtesy markers. Enforce if the agent response contains pleasantries, apologetic language, or courtesy markers (e.g. ’I’d be absolutely delighted to help!’, ’I’m so sorry, but…’, or ’Thank you so much for this wonderful question!’). brown1987politeness
@@
When the agent is solving a problem or explaining a concept, you prefer responses that state relevant assumptions, show step-by-step reasoning, and justify conclusions. Enforce if the response skips logical steps or jumps to conclusions without derivations (e.g. ’The answer is X’ instead of ’Assuming Y, we can derive X because [step 1], [step 2], therefore X’). When the agent is solving a problem or explaining a concept, you prefer responses that start with high-level intuition and generalizable principles before diving into the specific solution. Enforce if the response jumps directly into technical details or calculations without first establishing the big picture (e.g. starting with ’First, calculate X using formula Y…’ instead of ’The key principle here is Z, which applies whenever you see W. Let’s apply it to your problem…’). trope2012construal
@@
When working on a multi-step problem you prefer that each agent response covers only a single small increment of the problem, and asks for confirmation before proceeding to the next increment. Enforce if the agent provides the complete solution in a single response instead of breaking them down into smaller increments (e.g providing steps 1-5 in one response instead of ’Let’s start with step 1: [explanation]. Does this make sense before we continue?’). When working on a problem, you prefer holistic responses that address the full solution. Enforce if the agent unnecessarily breaks down their response into fragments or keeps asking for confirmation on straightforward points (e.g. ’Let me explain just part 1 first. [Brief explanation]. Should I continue?’ when you could handle the complete answer). kalyuga2009expertise
@@
You prefer the agent to always end responses with a proactive suggestion or next step. Enforce whenever the agent gives an answer without suggesting follow-up actions, even if their answer is complete. (e.g. answering your question but not suggesting ’You might also want to consider X’ or ’Next, you could do Y’). You prefer the agent to respond to only your request, and does not provide unsolicited suggestions or next steps. Enforce if the agent adds suggestions or next steps (e.g. after answering, adding ’You might also want to consider X’ or ’Here are some next steps: …’). horvitz1999principles
@@
You prefer the agent to offer key takeaways or a summary for future reference. Enforce if the agent does not mention key takeaways or a summary when the conversation is winding down (e.g. not suggesting ’Let me summarize the key points we covered’). chi1989self
@@
When given an answer, you prefer the agent to provide multiple viable approaches along with the tradeoffs for each. Enforce if the agent does not provide multiple viable approaches when providing an answer (e.g. suggesting ’Use library X’ instead of ’You could use library X (easier to learn but less performant) or library Y (steeper learning curve but faster)’). schwartz2002maximizing
@@
When the agent is providing an answer, you prefer the response to start out with an outline of what will be covered. Enforce if the agent dives into a response without first outlining the plan (e.g. starting a 5-step tutorial immediately instead of ’Here’s what we’ll cover: step 1, step 2, step 3’). scott1995decision
@@
When the agent provides an explanation, you prefer responses that include examples, analogies or metaphors. Enforce if explanations do not contain examples. sweller2006worked
@@
You prefer the agent responses to be no longer than three sentences. Enforce if the agent response exceeds three sentences. You prefer the agent responses structured in bullet-point format. Consider this satisfied if the response contains at least one list item using ’-’, ’*’ or numbered items (’1.’, ’2.’, …). Mixed prose + bullets is acceptable. Enforce only if the response contains no bullets at all. You prefer the agent structured responses that have numbered steps. Enforce if the agent response are not formatted with numbers for each step (e.g. ’1. X, 2. Y, 3. Z’ instead of ’X, Y, Z’). You prefer structured agent responses that have headings for each section. Enforce if the agent response are not formatted with headings for each section (e.g. ’X, ## Y, ### Z’ instead of ’X, Y, Z’). You prefer the agent responses that have a one-line TL;DR at the end. Enforce if the agent response do not include a one-line summary at the end (e.g. ’TL;DR: X’ instead of ’X’). When the agent provides an answer, you prefer the agent responses that contain confidence estimates. Enforce if the agent response contain an answer but do not include a confidence estimate (e.g. ’I’m 90% confident that X’, instead of ’X’). wood2007new
@@ Appendix C Environment Validation
  • Multi-Turn Interaction (S2): The problem statement is provided to a user simulator (with no interaction preferences) who engages in conversation with the agent. The agent must ask clarification questions to fully understand the problem statement, then provide an answer to the problem. This mirrors prior works such as wu2025collabllm, laban2025llms, zhou2025sweet. This setting isolates the effect of multi-turn interaction, introducing challenges such as underspecification and incorporating user input, and reveals how performance changes when agents must interact with users rather than solving problems directly.
@@ Discussion.
The transition from S1 to S2 reveals a notable drop in performance when moving from direct problem-solving to multi-turn interaction. This result aligns with findings from prior works (wu2025collabllm; laban2025llms; zhou2025sweet; mehri2025goal), and demonstrates how LLMs struggle with challenges in multi-turn interactions such as handling underspecified problem descriptions.
@@ Appendix D Training Hyperparameters
When training conversational agents to generate session-level reflections, we observed that models were prone to catastrophic forgetting and demonstrated degraded performance on other tasks such as response generation. We also observed that models generated excessively long responses, a known optimization bias in GRPO (liu2025understanding), and also compounded by the tendency of LLM judges to favor longer outputs (dubois2024length). To mitigate such issues and preserve general capabilities, careful hyperparameter selection was necessary.
We use LLaMa-Factory (zheng2024llamafactory) for SFT and VERL (sheng2025hybridflow) for GRPO. In Table 4, we summarize the hyperparameters that we used for both training stages. The hyperparameters that are not mentioned here are set to their default values.

### v2

@@ 9 Limitations
Our work provides a foundation for multi-session collaboration, with many promising directions for future research. We use Llama-3.3-70B-Instruct as our user simulator. While it is a highly capable model, user simulation requires reliably following complex instructions, such as consistently enforcing preferences throughout interactions, and our evaluation metrics depend on this behavior. We found that the simulator struggled to exhibit expected behaviors during more complex tasks such as coding, which we consequently excluded from our benchmark. Additionally, user simulators can be extended to capture more realistic human behaviors. For example, real users can express preferences implicitly through behavioral cues rather than explicitly. Or users can have more complex preferences, such as context-dependent preferences that vary across domains, or evolving preferences that change over time. Although we conduct a human user study to validate our findings across realistic tasks and real users, we acknowledge these limitations remain.
@@ Appendix A User Interaction Preferences
In human-AI collaborative settings, users often have specific interaction preferences that describe how they expect agents to behave. When agents adhere to these preferences, users may learn more effectively, communicate more efficiently, or experience smoother, more personalized interactions (Shi et al., 2025; Sweller, 1988). To ensure our benchmark captures realistic preferences, we draw inspiration from studies in psychology, cognitive science, and human-computer interaction. We manually curate these preferences to be well-suited for our collaborative problem-solving setting. Table LABEL:tab:interaction_preferences provides the complete taxonomy of interaction preferences used in the MultiSessionCollab benchmark.
@@
You prefer agent responses to avoid any unnecessary narration and get directly to the point. Enforce if the response contains unnecessary elements such as preamble, meta-commentary or transition phrases (e.g. ’That’s a great question!’ or ’Let me walk you through this…’). You prefer agent responses to be narrative, conversational, and engaging. The response should acknowledge previous context (e.g. ’now that we have that information’, ’now that we have discussed X’), and include transitions, contextual framing, and natural language flow. Enforce if the response is overly terse or choppy (e.g. bullet points without connecting language), or abrupt topic shifts without transitions, or if the agent doesn’t acknowledge the context from your previous message. Gudykunst et al. (1988)
@@
You prefer the agent responses to be respectful, considerate, and friendly. Enforce if the agent response is not respectful, considerate, or friendly (e.g. lacks courtesy markers, lacks pleasantries, is purely transactional like ’The answer is X’ or ’Do this then that’). You prefer the agent responses to be blunt without pleasantries, apologetic language, or courtesy markers. Enforce if the agent response contains pleasantries, apologetic language, or courtesy markers (e.g. ’I’d be absolutely delighted to help!’, ’I’m so sorry, but…’, or ’Thank you so much for this wonderful question!’). Brown (1987)
@@
When the agent is solving a problem or explaining a concept, you prefer responses that state relevant assumptions, show step-by-step reasoning, and justify conclusions. Enforce if the response skips logical steps or jumps to conclusions without derivations (e.g. ’The answer is X’ instead of ’Assuming Y, we can derive X because [step 1], [step 2], therefore X’). When the agent is solving a problem or explaining a concept, you prefer responses that start with high-level intuition and generalizable principles before diving into the specific solution. Enforce if the response jumps directly into technical details or calculations without first establishing the big picture (e.g. starting with ’First, calculate X using formula Y…’ instead of ’The key principle here is Z, which applies whenever you see W. Let’s apply it to your problem…’). Trope and Liberman (2012)
@@
When working on a multi-step problem you prefer that each agent response covers only a single small increment of the problem, and asks for confirmation before proceeding to the next increment. Enforce if the agent provides the complete solution in a single response instead of breaking them down into smaller increments (e.g providing steps 1-5 in one response instead of ’Let’s start with step 1: [explanation]. Does this make sense before we continue?’). When working on a problem, you prefer holistic responses that address the full solution. Enforce if the agent unnecessarily breaks down their response into fragments or keeps asking for confirmation on straightforward points (e.g. ’Let me explain just part 1 first. [Brief explanation]. Should I continue?’ when you could handle the complete answer). Kalyuga (2009)
@@
You prefer the agent to always end responses with a proactive suggestion or next step. Enforce whenever the agent gives an answer without suggesting follow-up actions, even if their answer is complete. (e.g. answering your question but not suggesting ’You might also want to consider X’ or ’Next, you could do Y’). You prefer the agent to respond to only your request, and does not provide unsolicited suggestions or next steps. Enforce if the agent adds suggestions or next steps (e.g. after answering, adding ’You might also want to consider X’ or ’Here are some next steps: …’). Horvitz (1999)
@@
You prefer the agent to offer key takeaways or a summary for future reference. Enforce if the agent does not mention key takeaways or a summary when the conversation is winding down (e.g. not suggesting ’Let me summarize the key points we covered’). Chi et al. (1989)
@@
When given an answer, you prefer the agent to provide multiple viable approaches along with the tradeoffs for each. Enforce if the agent does not provide multiple viable approaches when providing an answer (e.g. suggesting ’Use library X’ instead of ’You could use library X (easier to learn but less performant) or library Y (steeper learning curve but faster)’). Schwartz et al. (2002)
@@
When the agent is providing an answer, you prefer the response to start out with an outline of what will be covered. Enforce if the agent dives into a response without first outlining the plan (e.g. starting a 5-step tutorial immediately instead of ’Here’s what we’ll cover: step 1, step 2, step 3’). Scott and Bruce (1995)
@@
When the agent provides an explanation, you prefer responses that include examples, analogies or metaphors. Enforce if explanations do not contain examples. Sweller (2006)
@@
You prefer the agent responses to be no longer than three sentences. Enforce if the agent response exceeds three sentences. You prefer the agent responses structured in bullet-point format. Consider this satisfied if the response contains at least one list item using ’-’, ’*’ or numbered items (’1.’, ’2.’, …). Mixed prose + bullets is acceptable. Enforce only if the response contains no bullets at all. You prefer the agent structured responses that have numbered steps. Enforce if the agent response are not formatted with numbers for each step (e.g. ’1. X, 2. Y, 3. Z’ instead of ’X, Y, Z’). You prefer structured agent responses that have headings for each section. Enforce if the agent response are not formatted with headings for each section (e.g. ’X, ## Y, ### Z’ instead of ’X, Y, Z’). You prefer the agent responses that have a one-line TL;DR at the end. Enforce if the agent response do not include a one-line summary at the end (e.g. ’TL;DR: X’ instead of ’X’). When the agent provides an answer, you prefer the agent responses that contain confidence estimates. Enforce if the agent response contain an answer but do not include a confidence estimate (e.g. ’I’m 90% confident that X’, instead of ’X’). Wood and Neal (2007)
@@ Appendix B Agent Prompts
The effectiveness of our benchmark as an evaluation framework depends on user simulators that can reliably adhere to their user profiles and consistently enforce their preferences (Mehri et al., 2025). To ensure this, we employ a structured reasoning process that explicitly considers whether (a) each preference has been satisfied and if any needs to be enforced, (b) the draft answer should be updated, and (c) the conversation should be terminated. This structure yields interpretable user behavior signals: we track which utterances enforce preferences, using this both to quantify user effort (Section 3.3) and to derive learning signals for our RL framework (Section 4.2). The system prompt is provided in Appendix B.
@@ Appendix C Environment Validation
  • Multi-Turn Interaction (S2): The problem statement is provided to a user simulator (with no interaction preferences) who engages in conversation with the agent. The agent must ask clarification questions to fully understand the problem statement, then provide an answer to the problem. This mirrors prior works such as Wu et al. (2025b), Laban et al. (2025), Zhou et al. (2025a). This setting isolates the effect of multi-turn interaction, introducing challenges such as underspecification and incorporating user input, and reveals how performance changes when agents must interact with users rather than solving problems directly.
@@ Discussion.
The transition from S1 to S2 reveals a notable drop in performance when moving from direct problem-solving to multi-turn interaction. This result aligns with findings from prior works (Wu et al., 2025b; Laban et al., 2025; Zhou et al., 2025a; Mehri et al., 2025), and demonstrates how LLMs struggle with challenges in multi-turn interactions such as handling underspecified problem descriptions.
@@ Appendix D Training Hyperparameters
When training conversational agents to generate session-level reflections, we observed that models were prone to catastrophic forgetting and demonstrated degraded performance on other tasks such as response generation. We also observed that models generated excessively long responses, a known optimization bias in GRPO (Liu et al., 2025b), and also compounded by the tendency of LLM judges to favor longer outputs (Dubois et al., 2024). To mitigate such issues and preserve general capabilities, careful hyperparameter selection was necessary.
We use LLaMa-Factory (Zheng et al., 2024) for SFT and VERL (Sheng et al., 2025) for GRPO. In Table 5, we summarize the hyperparameters that we used for both training stages. The hyperparameters that are not mentioned here are set to their default values.